OpenStudy (anonymous):

Find Each segment length. Tell whether the segments are congruent AB: A(0,2), B(-3,8) CD: C(-2,2), D(0,-4)

OpenStudy (anonymous):

what is the formular?

OpenStudy (anonymous):

\[d=\sqrt{(x2-x1)^{2} + (y2-y1)^{2}}\]

OpenStudy (anonymous):

so to find AB you already know the coordinates so: \[AB= \sqrt{(-3-0)^{2}+(8-2)^{2}}\]

OpenStudy (anonymous):

\[AB=\sqrt{(-3)^{2} +5^{2}}\]

OpenStudy (anonymous):

\[AB=\sqrt{9+25}\]

OpenStudy (anonymous):

\[AB=\sqrt{34}\]

OpenStudy (anonymous):

You could leave AB = square roof 34 or round it to 5.83

OpenStudy (anonymous):

Do the same to CD: CD= spuare roof of [ (0- (-2) ^2) + (-4-2)^2 ] CD= square roof of [ ( 0+2)^2 + (-6)^2 ] CD= square roof of ( 4 +36) CD = square roof of 40 CD = 2 square roof 10 or 6.32

OpenStudy (anonymous):

congruents or not based on it definition was : exacly the same. Since you know 2 of it leagth, can you said it was congruents or not? 5.83 = 6.32 ( false ) => it NOT CONGRUENT! :)